OpenStudy (anonymous):

-6f+13=2f-11 please help :(

jimthompson5910 (jim_thompson5910):

Start off by subtracting 13 from both sides

jimthompson5910 (jim_thompson5910):

What do you get when you do this?

OpenStudy (anonymous):

what do you subtract 13 from?

jimthompson5910 (jim_thompson5910):

on the left side, 13-13 = 0 so that goes away on the left side

jimthompson5910 (jim_thompson5910):

on the right side -11 - 13 = -24

jimthompson5910 (jim_thompson5910):

so we go from this -6f+13=2f-11 to this -6f = 2f - 24

jimthompson5910 (jim_thompson5910):

what's next?

OpenStudy (anonymous):

i believe that you divide?

OpenStudy (anonymous):

sorry . if that's not correct. i just started learning this :(

jimthompson5910 (jim_thompson5910):

close, you actually subtract 2f from both sides

jimthompson5910 (jim_thompson5910):

you divide once you have all the variable terms on one side (and everything else on the other)

jimthompson5910 (jim_thompson5910):

doing that gives you -6f = 2f - 24 -6f = 2f - 24 -6f-2f = 2f - 24-2f -8f = -24

jimthompson5910 (jim_thompson5910):

Now you divide

OpenStudy (anonymous):

-6f=2f-24 -8f=-24 f=3

OpenStudy (anonymous):

remember - - =+

OpenStudy (anonymous):

thanks.